Determine whether the equation defines y as a function of x:
x-3y=9
This is a take home quizz....this particular problem confuses me.
Thank you so much for your help!!
Livey

Answer by ozymandias(9)

In that particular form, it does not directly look like it shows y as a function of x, but with a bit of jiggling around, we can move things about until we get :
x - 3y = 9
x = 9 + 3y
3y = x - 9
y = (x-9)/3 --> TA DAH !! y is a function of x !